# MoLoRec: A Generalizable and Efficient Framework for LLM-Based Recommendation

### v1

@@ 3.4. Discussion
Key Advantages of MoLoRec. 1) Generalization. MoLoRec is generalizable to various recommendation scenarios. For example, when facing a new recommendation domain, it only needs to train a new domain-specific LoRA module, enabling rapid generalization. Even in extreme cases where no training data is available for the new domain, MoLoRec can still work using the generalizable base model. For the new user or new item recommendation scenario, MoLoRec dynamically balances generalization and domain-specific specialization to deliver accurate recommendations. Additionally, leveraging the in-context learning capabilities of LLMs, MoLoRec naturally exhibits task generalization. E.g., it can generate explainable recommendation results. 2) Efficiency. The proposed MoLoRec paradigm is designed for ease of use, offering plug-and-play integration, where the domain-general module is trained once, and the domain-specific plugin is incorporated with minimal fine-tuning.
@@
Comparision to Existing Methods. Traditional sequential recommendation models (e.g., GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018)) typically rely on explicit item IDs for modeling, which restricts their generalization ability across new domains or platforms. To this end, transferable methods have been explored for universal sequence representation learning. For instance, VQ-Rec (Hou et al., 2023)) and UniSRec (Hou et al., 2022) leverage text to represent items and employ contrastive pre-training strategies on language models to enhance transferability. However, these pretraining methods cannot be directly applied to LLMs. With the rise of LLMs, their strong domain and task generalization capabilities present new opportunities for improving the generalization of recommender systems. We categorize these approaches as breadth-oriented methods. For example, P5 (Geng et al., 2022) designs prompts to unify five recommendation tasks, establishing a unified text-to-text recommendation paradigm. In contrast, depth-oriented methods focus on deeply aligning LLMs with specific domains. For instance, TallRec (Bao et al., 2023) employs LoRA for efficient domain adaptation. Another research direction explores aligning collaborative signals with LLMs, as seen in works like LLaRA (Liao et al., 2024) and iLoRA (Kong et al., 2024). While incorporating collaborative signals significantly enhances performance in warm-start scenarios, it comes at the cost of reduced generalization, making adaptation to new domains and cold-start situations more challenging.
4. Experiments
4.1. Experimental Settings
Methods Beauty Toys Sports MovieLens-1M NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 Traditional BPR-MF 0.1630 0.2588 0.1276 0.2056 0.1496 0.2338 0.1724 0.4185 GRU4Rec 0.1672 0.2752 0.1320 0.2243 0.1787 0.2829 0.1724 0.4423 SASRec 0.2410 0.3284 0.2223 0.3105 0.1957 0.2967 0.2257 0.4708 FMLP-Rec 0.2988 0.4000 0.2994 0.3990 0.2645 0.3812 0.2410 0.5515 Transferable UniSRec 0.2654 0.4089 0.2612 0.3998 0.2341 0.3721 0.2615 0.5594 VQ-Rec 0.2714 0.4157 0.2715 0.4119 0.2476 0.3944 0.2805 0.5745 LLM-Based Qwen2-7B 0.0300 0.0394 0.0843 0.1062 0.0170 0.0242 0.0814 0.1057 RecFormer 0.2858 0.3840 0.3001 0.3880 0.2667 0.3885 0.2743 0.5701 P5 0.1775 0.2482 0.1171 0.1709 0.1860 0.2674 0.2046 0.2947 TALLRec 0.3208 0.3479 0.3308 0.3583 0.3002 0.3274 0.4759 0.4971 Ours MoLoRec-G 0.3081 0.3316 0.2957 0.3209 0.2750 0.2998 0.5680 0.5918 MoLoRec-S 0.4079 0.4291 0.4076 0.4314 0.3735 0.3925 0.5460 0.5703 MoLoRec 0.4132* 0.4350* 0.4097* 0.4334* 0.3754* 0.3944* 0.5783* 0.6023*
Methods Beauty Toys Sports MovieLens-1M NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 Traditional BPR-MF 0.0306 0.0688 0.0333 0.0765 0.0350 0.0739 0.0723 0.1421 GRU4Rec 0.0562 0.1063 0.0447 0.0926 0.0640 0.0996 0.0798 0.1489 SASRec 0.0656 0.1368 0.0670 0.1210 0.0547 0.1203 0.0912 0.1891 FMLP-Rec 0.0587 0.1229 0.0537 0.1117 0.0545 0.1236 0.1145 0.1947 Transferable UniSRec 0.0957 0.1457 0.0814 0.1559 0.0832 0.1408 0.0985 0.1343 VQ-Rec 0.1189 0.1589 0.0957 0.1603 0.0985 0.1463 0.1025 0.1412 LLM-Based Qwen2-7B 0.0187 0.0260 0.0293 0.0356 0.0213 0.0273 0.0318 0.0407 RecFormer 0.1051 0.1687 0.0913 0.1592 0.0922 0.1489 0.1108 0.1547 P5 0.0871 0.1466 0.0755 0.1358 0.0758 0.1355 0.0957 0.1319 TALLRec 0.1415 0.1674 0.1251 0.1524 0.1226 0.1486 0.1458 0.1668 Ours MoLoRec-G 0.1746 0.2072 0.1474 0.1710 0.1581 0.1868 0.1455 0.1890 MoLoRec-S 0.1603 0.1863 0.1504 0.1764 0.1564 0.1867 0.1636 0.2597 MoLoRec 0.1825* 0.2145* 0.1614* 0.1821* 0.1646* 0.1921* 0.1818* 0.2755*
4.1.1. Datasets.
@@ 4.1. Experimental Settings
We conduct experiments on e-commerce and movie recommendation scenarios. For the e-commerce recommendation scenario, the domain-general instruction tuning dataset is conducted using seven e-commerce domains in Amazon111https://jmcauley.ucsd.edu/data/amazon/. and validated on three domain-specific datasets in Amazon (Beauty, Toys, Sports). For the movie recommendation scenario, the domain-general dataset is built using MovieLens-10M222https://grouplens.org/datasets/movielens/ and validated on the domain-specific dataset MovieLens-1M.
@@
For all datasets, items are represented using their textual ”title” information. To prevent data leakage, we carefully removed the overlapping portions between the domain-general dataset and the domain-specific datasets. We consider two recommendation settings: 1) Warm-Start Setting keeps the five-core dataset and filters users and items with fewer than five interactions for all datasets. Following (Geng et al., 2022; Lin et al., 2024b), we adopt the leave-one-out strategy to split the filtered dataset. More concretely, we split the last interaction of each user into the test set, the second-to-last one into the validation set, and the rest into the training data. 2) New-Item Setting uses the same training and validation sets as the warm-start setting, but replaces the items in the test set with those that never appear in the training or validation sets. (See Appendix A.1 for dataset statistics.)
4.1.2. Baselines.
@@
We compare MoLoRec with traditional recommendation methods (BPR-MF (Rendle et al., 2009), GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and FMLP-Rec) (Zhou et al., 2022), transferable sequential recommenders (UniSRec (Hou et al., 2022), VQ-Rec (Hou et al., 2023)), LLM-based recommenders (Qwen2-7B (Yang et al., 2024c), RecFormer (Li et al., 2023), P5 (Geng et al., 2022), TALLRec (Bao et al., 2023)) and our ablation counterparts (MoLoRec-G, MoLoRec-S). (See Appendix A.2 for more details of these baselines.)
4.1.3. Evaluation Setting.
@@
Following some previous LLM-based recommendation works (Zhang et al., 2024; Kim et al., 2024b), to evaluate the performance of the sequential recommendation models, we add 29 randomly selected non-interacted items to the test set, so that the test set of each user contains 1 positive item and 29 negative items. For quantitative comparison, we employ widely used ranking-based metrics, NDCG@1 and NDCG@3 for all experiments. All metrics are ”the higher, the better”. For all tables in the following, bold* numbers refer to the best performance, while underlined numbers indicate the second-best performance.
(See Appendix A.3 for implementation details.)
4.2. Overall Performance
@@
We comprehensively compare MoLoRec against traditional, transferable, and LLM-based recommenders. The experimental results on the warm-start I.I.D setting and the cold-start item out-of-distribution settings are shown in Table 1 and 2, respectively. From the experimental results, we have the following observations:
@@
The proposed MoLoRec consistently achieves the best performance across all I.I.D and O.O.D scenarios on the four datasets, with a t-test at p¡0.05 level. Specifically, in the warm-start scenario, MoLoRec achieves notable improvements in NDCG@1 over the best baseline methods (excluding our ablation counterparts), with performance gains of 28.8%, 23.85%, 25.05%, and 21.52% across the Beauty, Toys, Sports, and Movielens-1M, respectively. In the cold-start O.O.D scenario, the improvements are 28.98%, 29.02%, 34.26%, and 24.69%. The outstanding performance of MoLoRec in both I.I.D and O.O.D scenarios demonstrates its ability to efficiently capture domain-specific knowledge while exhibiting exceptional generalization capabilities.
Qwen2-7B demonstrates limited performance across all scenarios. However, the LLMs trained via TALLRec achieved significant improvements. This is because there is a gap between pre-training general text corpus of LLMs and the recommendation task, showing the importance of using recommendation knowledge for instruction fine-tuning on pre-trained LLMs.
@@
Traditional recommendation methods and the ID-based LLM recommendation method P5 perform poorly in cold-start scenarios. Relying heavily on collaborative filtering information reduces the model’s generalization capability.
4.3. In-Depth Analysis
4.3.1. Ablation Study
@@
From the performance of MoLoRec, MoLoRec-G, MoLoRec-S in the Table 1 and 2, we can find that: 1) MoLoRec consistently surpasses its ablation counterparts in all scenarios, which indicates that it is crucial for LLM to comprehend both general recommendation world knowledge and domain-specific knowledge. These two types of knowledge can complement each other. It also validates that the mixture-of-LoRA method can effectively integrate both types of knowledge. 2) Note that the MoLoRec-G has not been exposed to training data from the Beauty, Toys, Sports, or Movielens-1M domains in any experimental setting. However, it still achieves commendable performance in such zero-shot settings, demonstrating that it has effectively learned generalizable knowledge in the recommendation domain. This strong performance underscores its robust generalization ability.
4.3.2. Performance in Few-Shot Training Setting
We further conduct experiments in scenarios with limited domain-specific training data. Specifically, we adopt a few-shot training setup on MovieLens-1M, where only a small percentage of samples are randomly selected from the training set for model training. We compare MoLoRec with TallRec and the results of second-round fine-tuning on the generalizable base model. The experimental results are shown in Table 3. We find that the optimization approach of second-round fine-tuning led to catastrophic forgetting. It fails to generate output in the specified instruction format. Experimental results demonstrate that MoloRec maintains strong performance even in few-shot scenarios.
4.3.3. Analysis of the Coefficients and
@@
In Figure 4(a) and 4(b), we investigate the coefficients and calculated by entropy-guided adaptive mixture-of-LoRA in the warm-start scenario and the cold-start scenario, respectively. and represent the respective weights assigned to the domain-general and domain-specific LoRA module during their mixture. We observe that in the warm-start scenario, is relatively large, reflecting a greater reliance on domain-specific LoRA. Conversely, in the cold-start scenario, the weight of increases significantly, emphasizing the importance of domain-general knowledge. This result is reasonable and aligns with the differing requirements of these two scenarios. This result also demonstrates the effectiveness of our entropy-guided adaptive mixture-of-LoRA method.
4.3.4. Analysis of the Number of Unlabeled Test Data
@@
The number of unlabeled test data is one of the hyperparameters. Figure 5 illustrates the impact of different numbers on the NDCG@1 performance. The entropy-guided learning method converges rapidly, experimental results indicate that setting the number to 50 or 100 achieves a model fusion weight with optimal performance, this configuration proves effective across the majority of experiments conducted on the Beauty, Toys, Sports and MovieLens-1M datasets. Therefore, a limited amount of training data is sufficient to learn suitable model fusion weights, thereby significantly reducing the data requirements and computation resource costs.
4.3.5. Case Study
We further conduct a case study to delve deeper into the recommendation results of MoLoRec. We randomly selected a user from the Movielens-1M test set, provided their historical viewing records and a candidate set, and asked both our model and ChatGPT to make movie recommendations along with explanations for their choices. The outputs are shown in Figure 6. We find that MoLoRec successfully generalizes to the explainable recommendation task. It accurately captured the user’s preference for action movies from their historical viewing records and leveraged world knowledge to provide an accurate interpretation of the plot of Die Hard. In contrast, GPT-4 lacks domain-specific knowledge in the recommendation, incorrectly associating the action movie Die Hard with science fiction films like Star Wars, Alien, and Terminator, resulting in unreasonable explanation outcomes. This case study further highlights MoLoRec’s task generalization capability and its deep understanding of recommendation knowledge.
Scenario | Sample % | TallRec | 2nd Finetune | MoLoRec |
Warm-Start | 10% | 0.3957 | 0.0704 | 0.5353 |
@@ 4.3.5. Case Study
5. Related Works
Sequential Recommendation. Sequential recommendation utilizes users’ interaction histories to predict the next relevant item. Deep learning-based methods, such as RNNs (Hidasi et al., 2016; Li et al., 2017), GNNs (Qiu et al., 2019; Chang et al., 2021), and attention mechanisms (Kang and McAuley, 2018; Zhou et al., 2020; Xie et al., 2022), have become mainstream but rely solely on item IDs, limiting their adaptability to new scenarios. To improve transferability, transferable sequential recommendation studies (Hou et al., 2022, 2023; Li et al., 2023; Yuan et al., 2023) explore leveraging textual features to enhance item representations. These approaches improve the transferability and robustness of recommender systems. Nowadays, LLMs offer new opportunities for sequential modeling, promising more robust and generalizable recommender systems.
LLM-Based Recommendation. With the rise of LLMs, interest in LLM-based recommender systems has grown, leveraging LLMs as core engines. Early studies (Dai et al., 2023; Sanner et al., 2023; Wang et al., 2023) explore their zero-shot/few-shot potential via in-context learning (Dong et al., 2024). However, the gap between LLMs’ pretraining on general text and recommendation-specific needs leads to suboptimal performance. To address this, recent research follows two paradigms: breadth-oriented and depth-oriented. The former integrates multi-domain (Tang et al., 2024; Peng et al., 2024) or multi-task (Geng et al., 2022; Zhang et al., 2024; Cui et al., 2022; Peng et al., 2024) recommendation data to construct extensive recommendation world knowledge, paving the way for developing a generalizable LLM-based recommender. For example, Peng et al.(Peng et al., 2024) build large-scale e-commerce instruction dataset ECInstruct and develop generalist LLM for e-commerce. P5 (Geng et al., 2022) designs prompts to unify 5 recommendation tasks and presents a unified text-to-text recommendation paradigm. Depth-oriented paradigm seeks to enable LLMs to deeply comprehend recommendation tasks within specific domains. Key areas of focus include: 1) the in-depth alignment of domain-specific recommendation knowledge, such as collaborative signals (Lin et al., 2024b; Kim et al., 2024a; Liao et al., 2024; Ren et al., 2024; Kong et al., 2024). The introduction of collaborative signals effectively improves model performance in warm-start scenarios. However, this comes at the cost of reduced generalization, making it challenging to adapt to new domains and cold-start situations. Another research focus is 2) the development of efficient alignment methods between large language models and recommendation tasks. These methods include leveraging (Bao et al., 2023; Kong et al., 2024) LoRA fine-tuning techniques and designing data-efficient fine-tuning strategies (Lin et al., 2024c; Zheng et al., 2024). In fact, the two paradigms have complementary advantages. In this work, we investigate how to integrate the advantages of both paradigms to simultaneously achieve both breadth and depth.
Model Merging. Model merging aims to combine multiple expert models into a more capable single model, offering benefits such as reduced storage and serving costs (Yang et al., 2024a). Model merging techniques are applied in various scenarios such as unlearning old-knowledges in LLMs (Zaman et al., 2024), understanding content across multiple modalities (Aiello et al., 2024; Chen et al., 2024), generating images with different styles or achieving image-style transformation (Biggs et al., 2024). Previous attempts involve merging multiple models, all initially trained on the same task, with the aim of enhancing the model’s overall generalization (Gupta et al., 2020; Wang et al., 2022). Inspired by these works, we apply model merging to the LoRA modules, leveraging their ability to integrate domain-general and domain-specific knowledge effectively.
6. Conclusion
@@
In this paper, we proposed a generalizable and efficient LLM-based recommendation framework MoLoRec. As the recommendation data for a specific domain is limited, MoLoRec is designed to combine domain-general recommendation knowledge and domain-specific recommendation knowledge. Specifically, MoLoRec first aligned LLM with the general recommendation task by fine-tuning LLM from a multi-domain instruction dataset, and then modeled domain-specific recommendation tasks by fine-tuning the domain-specific instruction dataset. After that, the combination is achieved by merging two fine-tuning model parameters. MoLoRec is generalizable as it injects recommendation general knowledge to any domain-specific recommendation tasks. Besides, MoLoRec is efficient not only because it chooses parameter-efficient fine-tuning, but also the plug-and-play nature of any domain-specific recommendation task and domain-general task. Extensive experimental results on four recommendation datasets under both the warm scenario and cold-start scenario show the effectiveness of our proposed framework. E.g., MoLoRec shows about 30% average improvement on four datasets under cold-start scenarios.
Appendix A Appendix: Experimental Details

### v2

@@ 3.4 Discussion
Key Advantages of RecCocktail. 1) Generalization. RecCocktail is generalizable to various recommendation scenarios as it can adaptively determine how to fuse domain-general and domain-specific knowledge based on the distribution of the test data. Even in extreme cases where no training data is available for the new domain, RecCocktail can still work using the generalizable base spirit. Additionally, leveraging the in-context learning capabilities of LLMs, RecCocktail naturally exhibits task generalization. E.g., it can generate explainable recommendation results. 2) Efficiency. After adjusting the merging coefficients using a few unlabeled test data, we retains only a single LoRA module. As a result, there is no additional memory or computational overhead during inference. RecCocktail offers plug-and-play integration, where the domain-general module is trained once, and the domain-specific plugin is incorporated with minimal fine-tuning.
@@
Comparison to Existing Methods. Traditional sequential recommendation models (e.g., GRU4Rec (Hidasi et al. 2016), SASRec (Kang and McAuley 2018)) typically rely on explicit item IDs for modeling, limiting new domains or platforms generalization. Transferable methods address this by unifying multi-domain data in the input space. E.g., VQ-Rec (Hou et al. 2023) and UniSRec (Hou et al. 2022) use text representations with contrastive pre-training for enhanced transferability. With LLMs’ emergence, breadth-oriented methods leverage their generalization capabilities. P5 (Geng et al. 2022) unifies five recommendation tasks through a text-to-text paradigm. These methods enhance generalization by aggregating data from multiple domains/tasks. In contrast, RecCocktail takes a parameter merging approach, integrating knowledge directly in the parameter space—offering greater flexibility and scalability. Depth-oriented methods focus on domain-specific alignment. TallRec (Bao et al. 2023) uses LoRA for efficient adaptation, while LLaRA (Liao et al. 2024), iLoRA (Kong et al. 2024), and AlphaRec (Sheng et al. 2025) align collaborative signals with LLMs. Although this significantly enhances performance in warm-start scenarios, it reduces generalization for new domains and cold-start scenarios.
[…]
4 Experiments
4.1 Experimental Settings
@@
Datasets. We conduct experiments on e-commerce and movie recommendation scenarios. For the e-commerce recommendation scenario, the domain-general instruction tuning dataset is conducted using seven e-commerce domains in Amazon111https://jmcauley.ucsd.edu/data/amazon/. and validated on three domain-specific datasets in Amazon (Beauty, Toys, Sports). For the movie recommendation scenario, the domain-general dataset is built using MovieLens-10M222https://grouplens.org/datasets/movielens/ and validated on the domain-specific dataset MovieLens-1M.
@@
For all datasets, items are represented using their textual “title” information. To prevent data leakage, we carefully removed the overlapping portions between the domain-general dataset and the domain-specific datasets. We consider two recommendation settings: 1) Warm-Start Setting keeps the five-core dataset and filters users and items with fewer than five interactions for all datasets. Following (Geng et al. 2022; Lin et al. 2024a), we adopt the leave-one-out strategy to split the filtered dataset. More concretely, we split the last interaction of each user into the test set, the second-to-last one into the validation set, and the rest into the training data. 2) New-Item Setting uses the same training and validation sets as the warm-start setting, but replaces the items in the test set with those that never appear in the training or validation sets as ground-truth.
@@
Baselines. Baselines include traditional recommendation methods (BPR-MF (Rendle et al. 2009), GRU4Rec (Hidasi et al. 2016), SASRec (Kang and McAuley 2018), and FMLP-Rec) (Zhou et al. 2022), transferable sequential recommenders (UniSRec (Hou et al. 2022), VQ-Rec (Hou et al. 2023)), LLM-based recommenders (Qwen2-7B-zeroshot (Yang et al. 2024a), RecFormer (Li et al. 2023), P5 (Geng et al. 2022), TALLRec (Bao et al. 2023), AlphaRec (Sheng et al. 2025)) and our ablation counterparts (RecCocktail-WA, RecCocktail-G, RecCocktail-S).
@@
Evaluation Setting Following LLM-based recommendation works (Zhang et al. 2024; Kim et al. 2024b), we add 29 randomly selected non-interacted items to the candidate set, so that for each user it contains 1 positive item and 29 negative items. For quantitative comparison, we employ widely used ranking-based metrics, NDCG@1 and NDCG@3 for all experiments. All metrics are “the higher, the better”. For all tables in the following, bold* numbers refer to the best performance, while underlined numbers indicate the second-best performance.
(See Appendix A.1 for dataset statistics, Appendix A.2 for more baseline details, and Appendix A.3 for implementation details.)
4.2 Overall Performance
@@
We comprehensively compare RecCocktail against traditional, transferable, and LLM-based recommenders. Tables 1 and 2 present results for warm-start I.I.D and new-item O.O.D settings, respectively. Results indicate that:
@@
The proposed RecCocktail consistently achieves the best performance across all I.I.D and O.O.D scenarios on the four datasets, with a t-test at p0.05 level. The strong performance of RecCocktail demonstrates its ability to efficiently capture domain-specific knowledge while maintaining excellent generalization capabilities.
@@
Traditional recommendation methods and the ID-based LLM recommendation method P5, AlphaRec perform poorly in cold-start scenarios. Relying heavily on collaborative filtering information reduces the model’s generalization capability.
@@
We observe that RecCocktail consistently outperforms its ablation counterparts across all scenarios. This highlights the importance of integrating both general recommendation knowledge and domain-specific insights, which are complementary. The results also confirm the effectiveness of our LoRA cocktail method for knowledge fusion. Note that the RecCocktail-G has not been exposed to training data from the Beauty, Toys, Sports, or Movielens-1M domains. It still achieves commendable performance in such zero-shot settings, demonstrating that it has effectively learned generalizable recommendation knowledge. We also find that simple average merging (RecCocktail-WA) sometimes trigger negative transfer, demonstrating the effectiveness and importance of our proposed Entropy-Guided Adaptive Merging strategy for selecting appropriate merging coefficients.
4.3 In-Depth Analysis
@@
Analysis of the Coefficients and . In Figure 4(a) and 4(b), we investigate the coefficients and calculated by entropy-guided adaptive LoRA cocktail in the warm-start scenario and the cold-start scenario, respectively. and represent the respective weights assigned to the base spirit and the ingredient LoRA module during their merging. We observe that in the warm-start scenario, is relatively large, reflecting a greater reliance on domain-specific LoRA. Conversely, in the cold-start scenario, the weight of increases significantly, emphasizing the importance of domain-general knowledge. This result is reasonable and aligns with the differing requirements of these two scenarios. This result also demonstrates the effectiveness of our entropy-guided adaptive LoRA cocktail method.
@@
Analysis of the Number of Unlabeled Test Data. The number of unlabeled test data is one of the hyperparameters. Figure 4(c) and 4(d) illustrates the impact of different numbers on the NDCG@1 performance. The entropy-guided learning method converges rapidly, experimental results indicate that setting the number to 50 or 100 achieves a model fusion weight with optimal performance, this configuration proves effective across the majority of experiments conducted on the Beauty, Toys, Sports and MovieLens-1M datasets. Therefore, a limited amount of unlabeled test data is sufficient to learn suitable model fusion weights, thereby significantly reducing the data requirements and computation resource costs.
Performance in Few-Shot Training Setting.
| Scenario | Sample | TallRec | 2nd Finetune | RecCocktail |
| Warm | 10% | 0.3957 | 0.0704 | 0.5353 |
@@
We further conduct experiments in scenarios with limited domain-specific training data. Specifically, we adopt a few-shot training setup on MovieLens-1M, where only a small percentage of samples are randomly selected from the training set for model training. We compare RecCocktail with TallRec and the results of second-round fine-tuning on the generalizable base model. The experimental results are shown in Table 3. We find that the optimization approach of second-round fine-tuning led to catastrophic forgetting. It fails to generate output in the specified instruction format. The experimental results demonstrate that RecCocktail maintains strong performance even in few-shot scenarios.
(See Appendix A for more experimental results.)
5 Conclusion
@@
In this paper, we proposed a generalizable and efficient LLM-based recommendation framework RecCocktail. As the recommendation data for a specific domain is limited, RecCocktail is designed to combine domain-general and domain-specific recommendation knowledge. Specifically, RecCocktail first construct domain-general LoRA as base spirit and domain-specific LoRA as ingredient. Then the combination is achieved by merging two fine-tuning model parameters. RecCocktail is generalizable as it injects recommendation general knowledge to any domain-specific recommendation tasks. Besides, RecCocktail is efficient not only because it chooses parameter-efficient fine-tuning, but also the plug-and-play nature of any domain-specific recommendation task and domain-general task. Extensive experimental results on four recommendation datasets under both the warm scenario and cold-start scenario show the effectiveness of our proposed framework.
Appendix A Experimental Details